# TR-LLM: Integrating Trajectory Data for Scene-Aware LLM-Based Human Action Prediction

### v3

@@ II-B Language model-based approach
To the best of our knowledge, only a few studies have explored LLM-based human action prediction. [20] utilized an LLM to predict the next human activity by identifying the object in the environment that the person is most likely to interact with. [21] focused on LLM-based human action prediction to enable a robot to plan corresponding supportive actions. In their approach, they provided the LLM with object lists and action history to predict the next human action. However, while they account for scene context, they do not consider physical constraints such as human motion or the spatial structure of the environment.
Furthermore, vision-language models (VLMs)[34] have recently emerged as a prominent technology. These models possess the capability to recognize and understand both images and text, and they are utilized in a wide range of applications. Although they could potentially be applied to our task, pre-trained VLMs generally lack the ability to predict physical quantities under complex image constraints since they are primarily trained to establish correspondences between images and text. In contrast, our method can evaluate complex image constraints and predict the target object by employing a physically-aware prediction module that operates independently of the LLM module.
@@ IV-C Experimental setup
We provided three types of text-based scene context as input data to assess the model’s robustness under degraded conditions: (1) full scene context, (2) scene context without conversation, and (3) scene context without conversation and past action history. These variations simulate real-world scenarios where certain information may be unavailable. The evaluation examines how our method mitigates performance degradation, demonstrating its robustness to noise in practical settings. Additionally, we evaluate performance under three trajectory progress conditions: d¿1m, d¿2m, and d¿3m. The performance differences across these conditions reveal how the trajectory information contributes in each method.

### v4

@@ II-B Language model-based approach
Recently, several studies have explored LLM-based human action prediction. Specifically, [20, 21, 35] used LLMs to predict human actions, enabling robots to plan supportive responses. In these approaches, scene context was provided as textual input to the LLM for predicting subsequent human actions. However, while these studies incorporated contextual information, they neglected physical constraints such as realistic human motion dynamics or spatial environment structures. Meanwhile, [39] addressed human motion by introducing a graph-based transition model, but their representation of human movements was heuristic and simplistic, lacking grounding in real-world human trajectory data.
Furthermore, vision-language models (VLMs)[34] have recently emerged as a prominent technology. These models possess the capability to recognize and understand both images and text, and they are utilized in a wide range of applications such as action prediction[38] or spatial reasoning[36, 37]. Although they could potentially be applied to our task, pre-trained VLMs generally lack the ability to predict physical quantities under complex image constraints since they are primarily trained to establish correspondences between images and text. In contrast, our method can evaluate complex image constraints and predict the target object by employing a physically-aware prediction module that operates independently of the LLM module.
@@ IV-C Experimental setup
We provided three types of text-based scene context as input data to assess the model’s robustness under degraded conditions: (1) full scene context, (2) scene context without conversation, and (3) scene context without conversation and past action history. These variations simulate real-world scenarios where certain information may be unavailable. The evaluation examines how our method mitigates performance degradation, demonstrating its robustness to noise in practical settings. Additionally, we evaluate performance under three conditions of trajectory progress (d): $d>1\,\text{m}$, $d>2\,\text{m}$, and $d>3\,\text{m}$. The performance differences across these conditions reveal how the trajectory information contributes in each method.